4. Which has the higher annual yield, 6% compounded quarterly or 6.25% compounded semiannually?


1
Expert's answer
2022-06-09T11:01:03-0400
"A=P(1+r\/n)^{nt}"

6% compounded quarterly


"A=P(1+0.06\/4)^{4(1)}=1.061364P"

6.25% compounded semiannually


"A=P(1+0.0625\/2)^{2(1)}=1.063477P"

6.25% compounded semiannually has the higher annual yield.